OpenStudy (anonymous):

Simplify -(6p + 12) using the Distributive Property.

OpenStudy (anonymous):

-(6p + 12) = -1(6p) + -1(12) = ?

mathslover (mathslover):

K first of all @YunqVersatile_ Welcome to Openstudy : Distributive property: \(\huge{a(b+c)=ab+ac}\) Apply the same here also : what you have ? \(\huge{-1(6p+12)}\)

mathslover (mathslover):

Now according tot distributive property : \ \[\huge{-1(6p)-1(12)}\] I think u can do this further on ur own ? can u ?

OpenStudy (anonymous):

yes i can thanks.

mathslover (mathslover):

Ur welcome and best of luck for ur journey here in Openstudy